# Bigraded path homology and the magnitude-path spectral sequence

### 27 Mar 2026 v2

@@ 2.1. The category of directed graphs
2.2. Reachability chains and the length filtration
@@
Given a simplicial set $A$, we write $N(A)$ or simply $NA$ for the complex of normalized simplicial chains on $A$; if $A$ is filtered, we equip $NA$ with the filtration defined by $F_{\ell}NA=N(F_{\ell}A)$. The reachability complex of $G$ is precisely the normalized complex of simplicial chains in the nerve of $G$ [26, Section 4]
@@ 2.4. The $E^{1}$-page
  • Regard the cycle as an element of $\mathrm{RC}_{i+j}(G)$. It is a combination of generators of length $i$.
@@ 2.5. The $E^{2}$-page and beyond
Thus, for example, there is a $1$-homotopy from $f$ to $g$ if, for each $x$, either $f(x)$ and $g(x)$ are equal or there is a directed edge from the former to the latter. In general, $r$-homotopy (or long homotopy) is a condition on the pair of maps $f$ and $g$ which requires the existence of certain paths in $H$, but does not demand us to make a particular choice of such paths. The relation $\rightsquigarrow_{r}$ is not symmetric; nor, when $r\neq 0,\infty$, is it transitive. However, Asao proved the following, which we state explicitly here for future reference.
@@
In particular, this says that the entirety of the $E^{2}$-term of the MPSS is invariant under $1$-homotopy. (See also [4, Proposition 5.7] for this statement.) Meanwhile, reachability homology is invariant under long homotopy:
@@ 3. Bigraded path homology
We saw in the last section that the $E^{2}$-page of the MPSS of a directed graph $G$ contains the path homology $\mathrm{PH}_{\ast}(G)$ as its horizontal axis, and that the entire page has the same homotopy invariance property that $\mathrm{PH}_{\ast}(G)$ does, namely invariance under $1$-homotopy. This motivates the following definition.
@@
Let $X$ be a directed graph and let $A$ be a subgraph of $X$. Then $\mathrm{RC}(A)$ is a subcomplex of $\mathrm{RC}(X)$, and we define the relative reachability chains of the pair $(X,A)$ to be the quotient chain complex
@@
Recall that a subgraph $A$ of $X$ is said to be convex if for every pair of vertices $a,a^{\prime}$ in $A$ we have $d_{A}(a,a^{\prime})=d_{X}(a,a^{\prime})$ [34, Definition 4.2].
@@
Let $X$ be a graph and let $A$ be a subgraph of $X$. If $A$ is convex in $X$, then there is a short exact sequence of magnitude chains:
@@
If, in addition, there are no edges from $X\setminus A$ into $A$, then (3) is split by a chain map, (4) splits into short exact sequences, and we obtain a long exact sequence of bigraded path homology groups:
@@
Let us assume for the rest of the proof that there are no edges into $A$ from $X\setminus A$. This ‘no-entry’ condition on $A$ means that the only vertices of $X$ that admit paths into $A$ are those that already lie in $A$. (In particular, the only paths in $X$ between vertices of $A$ are those that lie wholly in $A$, so this condition alone ensures that $A\subseteq X$ is convex.)
We now show that (3) is split. Observe that, by the no-entry condition, a generator of $\mathrm{MC}(X)$ lies in $\mathrm{MC}(A)$ if and only if its final entry lies in $A$. Thus we may define a map $p\colon\mathrm{MC}(X)\to\mathrm{MC}(A)$ by the following rule.
@@ 4.1. Pairings and tensor products of spectral sequences
Let $A$ and $B$ be filtered chain complexes. Their (filtered) tensor product is the chain complex $A\otimes B$ equipped with the filtration in which
@@ 4.2. The Eilenberg–Zilber theorems
Here $r=k+k^{\prime}$, and $\sigma$ runs over all sequences $((i_{0},j_{0}),\ldots,(i_{r},j_{r}))$ in which $0\leq i_{s}\leq k$, $0\leq j_{s}\leq k^{\prime}$, and in which each term $(i_{s+1},j_{s+1})$ is obtained from $(i_{s},j_{s})$ by increasing exactly one of the components by $1$. The coefficient $\operatorname{sign}(\sigma)$ is defined to be $(-1)^{n}$ where $n$ is the number of pairs $(i,j)$ for which $i=i_{k}\implies j<j_{k}$.
@@
To prove Theorem 4.7, we will make use of a filtered variant of the classical Eilenberg–Zilber theorem. The classical theorem says that we have, for any simplicial sets $A$ and $B$, the Eilenberg–Zilber map $\nabla\colon NA\otimes NB\to N(A\times B)$ and the Alexander–Whitney map $\Delta\colon N(A\times B)\to NA\otimes NB$, which satisfy $\Delta\circ\nabla=\mathrm{Id}_{NA\otimes NB}$, and a chain homotopy $\mathrm{SHI}$ between $\nabla\circ\Delta$ and $\mathrm{Id}_{N(A\times B)}$. (For details, see Section 5 of [16] or Section 2 of [17]; in particular the latter reference contains explicit descriptions of $\Delta$, $\nabla$ and $\mathrm{SHI}$.) For present purposes, it is important to know that if $A$ and $B$ are filtered, then all these maps are automatically filtration-preserving.
@@
Let $A$ and $B$ be filtered simplicial sets. We equip their product $A\times B$ with the filtration given by $F_{\ell}(A\times B)=\bigcup_{i+j=\ell}F_{i}A\times F_{j}B$. Then the Alexander–Whitney map $\Delta\colon N(A\times B)\to NA\otimes NB$, the Eilenberg–Zilber map $\nabla\colon NA\otimes NB\to N(A\times B)$, and the chain homotopy $\mathrm{SHI}$ between $\nabla\circ\Delta$ and $\mathrm{Id}_{N(A\times B)}$ are all filtration-preserving. Thus, $NA\otimes NB$ and $N(A\times B)$ are chain homotopy equivalent as filtered chain complexes.
@@ 6.1. Cofibrations
Let $X$ be a directed graph, and $A\subseteq X$ a subgraph. The reach of $A$, denoted $rA$, is the induced subgraph of $X$ on the set of all vertices that admit a path from some vertex in $A$.
@@
    There are no edges from vertices not in $A$ to vertices in $A$.
@@
Before stating the theorems, we make a few remarks on the definition. First, note that condition (1) is equivalent to saying that there are no paths in $X$ from vertices outside $A$ to vertices inside $A$. This guarantees, in particular, that $A$ is a convex subgraph of $X$. Condition (2) says that $X$ projects to $A$ in the sense of Leinster [34, Definition 4.6] and its precursor (also Leinster) [33, Definition 2.3.1]. As noted in [34], the vertex $\pi(x)$ is the vertex of $A$ closest to $x$, and this determines $\pi(x)$ uniquely. In particular, taking $x=a$ in condition (2) we see that $\pi(a)=a$ for every $a\in V(A)$. Thus, we have a projection function $\pi\colon V(rA)\to V(A)$. In general, though, $\pi$ does not determine a map of graphs.
@@
Let $I$ denote the following directed graph.
@@
Given an arbitrary directed graph $X$, we define $CX$ to be the directed graph obtained from $X\square I$ by identifying the induced subgraph $X\square\{+1\}$ to a single vertex that we denote simply +1+1. We think of $CX$ as a form of ‘cone’ on $X$. We now identify $X$ with the induced subgraph of $CX$ on the vertices of form (x,1)(x,-1) for $x$ a vertex of $X$, so that we obtain the induced subgraph inclusion
@@
    Since there is no edge of $I$ from 0 to 1-1, there are no edges of $CX$ from $CX\setminus X$ to $X$.
@@
    The only paths in $I$ that begin at 1-1 are the trivial path and the edge $1\to 0$, so that the reach $rX$ of $X$ is the induced subgraph on the vertices of the form $(x,0)$ and (x,1)(x,-1) for $x$ a vertex of $X$. We then define $\pi$ on such vertices by
@@
Note also that $CX$ is 1-contractible: the identity map $CX\to CX$ is $1$-homotopic to the constant map with value +1+1. Indeed, if we denote by $d\colon CX\to CX$ the map defined by $d(+1)=+1$ and $d(x,j)=(x,\max\{j,0\})$ for $x\in X$ and j=0,1j=0,-1, then one can check that there are $1$-homotopies from both the identity map, and the constant map, to $d$. This example will be used later in Corollary 6.10.
@@
This superficial change means that Definition 6.2 can also be seen as a strengthening, suited to this paper’s filtered techniques, of the notion of long cofibration of directed graphs given in [26, Definition 6.2]: every cofibration in the sense of this paper is also a long cofibration. As is shown in [26, Proposition 6.6], a long cofibration $A\hookrightarrow X$ is precisely a map of directed graphs that induces a Dwyer morphism between the preorders generated by $A$ and $X$. The class of Dwyer morphisms contains all cofibrations in the Thomason model structure on the category of small categories, which is Quillen equivalent to the classical model structure on the category of simplicial sets [41, 12]. Thus, the theory being developed here and in Section 7 bears a close relationship to more classical homotopy theoretic constructions. The details of that relationship remain to be explored.
@@ 6.2. The excision and Mayer–Vietoris theorems
In the classical topological setting, the excision theorem for spaces gives rise to the Mayer–Vietoris sequence by ‘stitching together’, using a standard argument of homological algebra, the long exact sequences of pairs of the form $(X,A)$ and $(X\cup_{A}Y,Y)$. The same reasoning can be applied here to either the short exact sequences of magnitude homology groups, or the long exact sequences of bigraded path homology groups, that are associated to a pair by Theorem 3.3. That is how we obtain the next two theorems, which give a short exact Mayer–Vietoris sequence for magnitude homology, and a long exact Mayer–Vietoris sequence for path homology.
However, we are not at this time able to give a Mayer–Vietoris result for the $E^{r}$-pages for any $r\geq 3$. Indeed, the snake lemma allows us to pass from short exact sequences of chain complexes to long exact sequences of homology groups, and this is how we pass from $E^{1}$ to $E^{2}$ in the proof of Theorem 3.3. But we do not know of any account of the structure obtained from long exact sequences of chain complexes by taking homology, and this prevents us from obtaining results of Mayer–Vietoris type on the later pages of the MPSS.
@@
The category of undirected graphs embeds canonically into $\mathbf{DiGraph}$ by taking a graph $G$ and replacing each of its edges $\{x,y\}$ by a pair of directed edges $x\rightleftarrows y$. However, given a projecting decomposition $G=X\cup_{A}Y$, it is almost never the case that the inclusion of $A$ into $X$ (or into $Y$) induces, under this operation, a cofibration in the sense of Definition 6.2. Indeed, that occurs only when $X$ (respectively, $Y$) is the disjoint union of $A$ and its complement, in which case the split exact sequence (16) holds trivially for $\mathrm{MH}_{\ast,\ast}(X\cup_{A}Y)$. On the other hand, there are many non-trivial examples of projecting decompositions of undirected graphs; see, for example, Corollaries 4.13 and 4.14 of Leinster [34]. Thus, our Mayer–Vietoris theorem and that in [27] are quite independent.
@@
Since the inclusion $A\subseteq X$ satisfies the condition that there are no edges from $X\setminus A$ to $A$, Theorem 3.3 shows that $i_{\ast}$ is split, and it follows that $(i_{\ast},-f_{\ast})$ is also split, so that we obtain the short exact sequence of the claim. ∎
Proof of Theorem 6.8, assuming Theorem 6.6.
This is similar to the proof of Theorem 6.6, except that one uses the long exact sequences of bigraded path homology groups obtained from Theorem 3.3. We leave the details to the reader. ∎
@@
Let $X$ be a directed graph. We define $SX$ to be the graph obtained from $X$ by adding two new vertices, +1+1 and 1-1, together with an edge from each of the new vertices to each vertex of $X$. Thus $SX$ is a form of ‘unreduced suspension’ of $X$; see Example 6.11 for an illustrative example.
@@
Recall from Example 6.3 the graph $I$ given by $1\to 0\leftarrow+1$, and the graph $CX$ obtained from $X\square I$ by identifying $X\square\{+1\}$ to a single vertex +1+1. We showed that the inclusion $X\hookrightarrow CX$ that identifies $X$ with the subgraph $X\square\{-1\}$ is a cofibration. Now define a map $CX\to SX$ sending $X\square\{0\}$ to $X$, $X\square\{+1\}$ to +1+1, and $X\square\{-1\}$ to 1-1. Then we have a pushout square
@@
as one sees using the fact that $CX\to\bullet$ induces an isomorphism in path homology thanks to $1$-contractibility of $CX$. The three isomorphisms obtained in this paragraph are natural in $X$ and, combined, they complete the proof. ∎

### 27 Apr 2026 v3

@@ 2.1. The category of directed graphs
$2.$ Reachability chains and the length filtration
@@
Given a simplicial set AA, we write $N(A)$ or simply NANA for the complex of normalized simplicial chains on AA; if AA is filtered, we equip NANA with the filtration defined by $F_{\ell}NA=N(F_{\ell}A)$. The reachability complex of $G$ is precisely the normalized complex of simplicial chains in the nerve of $G$ [26, Section 4]
@@ 2.4. The $E^{1}$-page
  • Regard the cycle as an element of $\mathrm{RC}_{i+j}(G)$. It is a combination of generators of length ii.
@@ 2.5. The $E^{2}$-page and beyond
Thus, for example, there is a 11-homotopy from $f$ to $g$ if, for each $x$, either $f(x)$ and $g(x)$ are equal or there is a directed edge from the former to the latter. In general, $r$-homotopy (or long homotopy) is a condition on the pair of maps $f$ and $g$ which requires the existence of certain paths in $H$, but does not demand us to make a particular choice of such paths. The relation $\rightsquigarrow_{r}$ is not symmetric; nor, when $r\neq 0,\infty$, is it transitive. However, Asao proved the following, which we state explicitly here for future reference.
@@
In particular, this says that the entirety of the $E^{2}$-term of the MPSS is invariant under 11-homotopy. (See also [4, Proposition 5.7] for this statement.) Meanwhile, reachability homology is invariant under long homotopy:
@@ 3. Bigraded path homology
We saw in the last section that the $E^{2}$-page of the MPSS of a directed graph $G$ contains the path homology $\mathrm{PH}_{\ast}(G)$ as its horizontal axis, and that the entire page has the same homotopy invariance property that $\mathrm{PH}_{\ast}(G)$ does, namely invariance under 11-homotopy. This motivates the following definition.
@@
Let $X$ be a directed graph and let AA be a subgraph of $X$. Then $\mathrm{RC}(A)$ is a subcomplex of $\mathrm{RC}(X)$, and we define the relative reachability chains of the pair $(X,A)$ to be the quotient chain complex
@@
Recall that a subgraph AA of $X$ is said to be convex if for every pair of vertices $a,a^{\prime}$ in AA we have $d_{A}(a,a^{\prime})=d_{X}(a,a^{\prime})$ [34, Definition 4.2].
@@
Let $X$ be a graph and let AA be a subgraph of $X$. If AA is convex in $X$, then there is a short exact sequence of magnitude chains:
@@
If, in addition, there are no edges from $X\setminus A$ into AA, then (3) is split by a chain map, (4) splits into short exact sequences, and we obtain a long exact sequence of bigraded path homology groups:
@@
Let us assume for the rest of the proof that there are no edges into AA from $X\setminus A$. This ‘no-entry’ condition on AA means that the only vertices of $X$ that admit paths into AA are those that already lie in AA. (In particular, the only paths in $X$ between vertices of AA are those that lie wholly in AA, so this condition alone ensures that $A\subseteq X$ is convex.)
We now show that (3) is split. Observe that, by the no-entry condition, a generator of $\mathrm{MC}(X)$ lies in $\mathrm{MC}(A)$ if and only if its final entry lies in AA. Thus we may define a map $p\colon\mathrm{MC}(X)\to\mathrm{MC}(A)$ by the following rule.
@@ 4.1. Pairings and tensor products of spectral sequences
Let AA and $B$ be filtered chain complexes. Their (filtered) tensor product is the chain complex $A\otimes B$ equipped with the filtration in which
@@ 4.2. The Eilenberg–Zilber theorems
Here $r=k+k^{\prime}$, and $\sigma$ runs over all sequences $((i_{0},j_{0}),\ldots,(i_{r},j_{r}))$ in which $0\leq i_{s}\leq k$, $0\leq j_{s}\leq k^{\prime}$, and in which each term $(i_{s+1},j_{s+1})$ is obtained from $(i_{s},j_{s})$ by increasing exactly one of the components by 11. The coefficient $\operatorname{sign}(\sigma)$ is defined to be $(-1)^{n}$ where $n$ is the number of pairs $(i,j)$ for which $i=i_{k}\implies j<j_{k}$.
@@
To prove Theorem 4.7, we will make use of a filtered variant of the classical Eilenberg–Zilber theorem. The classical theorem says that we have, for any simplicial sets AA and $B$, the Eilenberg–Zilber map $\nabla\colon NA\otimes NB\to N(A\times B)$ and the Alexander–Whitney map $\Delta\colon N(A\times B)\to NA\otimes NB$, which satisfy $\Delta\circ\nabla=\mathrm{Id}_{NA\otimes NB}$, and a chain homotopy $\mathrm{SHI}$ between $\nabla\circ\Delta$ and $\mathrm{Id}_{N(A\times B)}$. (For details, see Section 5 of [16] or Section 2 of [17]; in particular the latter reference contains explicit descriptions of $\Delta$, $\nabla$ and $\mathrm{SHI}$.) For present purposes, it is important to know that if AA and $B$ are filtered, then all these maps are automatically filtration-preserving.
@@
Let AA and $B$ be filtered simplicial sets. We equip their product $A\times B$ with the filtration given by $F_{\ell}(A\times B)=\bigcup_{i+j=\ell}F_{i}A\times F_{j}B$. Then the Alexander–Whitney map $\Delta\colon N(A\times B)\to NA\otimes NB$, the Eilenberg–Zilber map $\nabla\colon NA\otimes NB\to N(A\times B)$, and the chain homotopy $\mathrm{SHI}$ between $\nabla\circ\Delta$ and $\mathrm{Id}_{N(A\times B)}$ are all filtration-preserving. Thus, $NA\otimes NB$ and $N(A\times B)$ are chain homotopy equivalent as filtered chain complexes.
@@ 6.1. Cofibrations
Let $X$ be a directed graph, and $A\subseteq X$ a subgraph. The reach of AA, denoted rArA, is the induced subgraph of $X$ on the set of all vertices that admit a path from some vertex in AA.
@@
    There are no edges from vertices not in AA to vertices in AA.
@@
Before stating the theorems, we make a few remarks on the definition. First, note that condition (1) is equivalent to saying that there are no paths in $X$ from vertices outside AA to vertices inside AA. This guarantees, in particular, that AA is a convex subgraph of $X$. Condition (2) says that $X$ projects to AA in the sense of Leinster [34, Definition 4.6] and its precursor (also Leinster) [33, Definition 2.3.1]. As noted in [34], the vertex $\pi(x)$ is the vertex of AA closest to $x$, and this determines $\pi(x)$ uniquely. In particular, taking $x=a$ in condition (2) we see that $\pi(a)=a$ for every $a\in V(A)$. Thus, we have a projection function $\pi\colon V(rA)\to V(A)$. In general, though, $\pi$ does not determine a map of graphs.
@@
Let II denote the following directed graph.
@@
Given an arbitrary directed graph $X$, we define CXCX to be the directed graph obtained from $X\square I$ by identifying the induced subgraph $X\square\{+1\}$ to a single vertex that we denote simply +1+1. We think of CXCX as a form of ‘cone’ on $X$. We now identify $X$ with the induced subgraph of CXCX on the vertices of form (x,1)(x,-1) for $x$ a vertex of $X$, so that we obtain the induced subgraph inclusion
@@
    Since there is no edge of II from 0 to 1-1, there are no edges of CXCX from $CX\setminus X$ to $X$.
@@
    The only paths in II that begin at 1-1 are the trivial path and the edge $1\to 0$, so that the reach rXrX of $X$ is the induced subgraph on the vertices of the form $(x,0)$ and (x,1)(x,-1) for $x$ a vertex of $X$. We then define $\pi$ on such vertices by
@@
Note also that CXCX is 1-contractible: the identity map $CX\to CX$ is 11-homotopic to the constant map with value +1+1. Indeed, if we denote by $d\colon CX\to CX$ the map defined by $d(+1)=+1$ and $d(x,j)=(x,\max\{j,0\})$ for $x\in X$ and j=0,1j=0,-1, then one can check that there are 11-homotopies from both the identity map, and the constant map, to $d$. This example will be used later in Corollary 6.10.
@@
This superficial change means that Definition 6.2 can also be seen as a strengthening, suited to this paper’s filtered techniques, of the notion of long cofibration of directed graphs given in [26, Definition 6.2]: every cofibration in the sense of this paper is also a long cofibration. As is shown in [26, Proposition 6.6], a long cofibration $A\hookrightarrow X$ is precisely a map of directed graphs that induces a Dwyer morphism between the preorders generated by AA and $X$. The class of Dwyer morphisms contains all cofibrations in the Thomason model structure on the category of small categories, which is Quillen equivalent to the classical model structure on the category of simplicial sets [41, 12]. Thus, the theory being developed here and in Section 7 bears a close relationship to more classical homotopy theoretic constructions. The details of that relationship remain to be explored.
@@ 6.2. The excision and Mayer–Vietoris theorems
In the classical topological setting, the excision theorem for spaces gives rise to the Mayer–Vietoris sequence by ‘stitching together’, using a standard argument of homological algebra, the long exact sequences of pairs of the form $(X,A)$ and $(X\cup_{A}Y,Y)$. The same reasoning can be applied here to either the short exact sequences of magnitude homology groups, or the long exact sequences of bigraded path homology groups, that are associated to a pair by Theorem $3.$ That is how we obtain the next two theorems, which give a short exact Mayer–Vietoris sequence for magnitude homology, and a long exact Mayer–Vietoris sequence for path homology.
However, we are not at this time able to give a Mayer–Vietoris result for the $E^{r}$-pages for any $r\geq 3$. Indeed, the snake lemma allows us to pass from short exact sequences of chain complexes to long exact sequences of homology groups, and this is how we pass from $E^{1}$ to $E^{2}$ in the proof of Theorem $3.$ But we do not know of any account of the structure obtained from long exact sequences of chain complexes by taking homology, and this prevents us from obtaining results of Mayer–Vietoris type on the later pages of the MPSS.
@@
The category of undirected graphs embeds canonically into $\mathbf{DiGraph}$ by taking a graph $G$ and replacing each of its edges $\{x,y\}$ by a pair of directed edges $x\rightleftarrows y$. However, given a projecting decomposition $G=X\cup_{A}Y$, it is almost never the case that the inclusion of AA into $X$ (or into $Y$) induces, under this operation, a cofibration in the sense of Definition 6.2. Indeed, that occurs only when $X$ (respectively, $Y$) is the disjoint union of AA and its complement, in which case the split exact sequence (16) holds trivially for $\mathrm{MH}_{\ast,\ast}(X\cup_{A}Y)$. On the other hand, there are many non-trivial examples of projecting decompositions of undirected graphs; see, for example, Corollaries 4.13 and 4.14 of Leinster [34]. Thus, our Mayer–Vietoris theorem and that in [27] are quite independent.
@@
Since the inclusion $A\subseteq X$ satisfies the condition that there are no edges from $X\setminus A$ to AA, Theorem 3.3 shows that $i_{\ast}$ is split, and it follows that $(i_{\ast},-f_{\ast})$ is also split, so that we obtain the short exact sequence of the claim. ∎
Proof of Theorem 6.8, assuming Theorem $6.$
This is similar to the proof of Theorem 6.6, except that one uses the long exact sequences of bigraded path homology groups obtained from Theorem $3.$ We leave the details to the reader. ∎
@@
Let $X$ be a directed graph. We define SXSX to be the graph obtained from $X$ by adding two new vertices, +1+1 and 1-1, together with an edge from each of the new vertices to each vertex of $X$. Thus SXSX is a form of ‘unreduced suspension’ of $X$; see Example 6.11 for an illustrative example.
@@
Recall from Example 6.3 the graph II given by $1\to 0\leftarrow+1$, and the graph CXCX obtained from $X\square I$ by identifying $X\square\{+1\}$ to a single vertex +1+1. We showed that the inclusion $X\hookrightarrow CX$ that identifies $X$ with the subgraph $X\square\{-1\}$ is a cofibration. Now define a map $CX\to SX$ sending $X\square\{0\}$ to $X$, $X\square\{+1\}$ to +1+1, and $X\square\{-1\}$ to 1-1. Then we have a pushout square
@@
as one sees using the fact that $CX\to\bullet$ induces an isomorphism in path homology thanks to 11-contractibility of CXCX. The three isomorphisms obtained in this paragraph are natural in $X$ and, combined, they complete the proof. ∎